OpenStudy (anonymous):

Find the circumference and area of a circle with a diameter of 22 inches. Leave your answers in terms of pi. Select one: a. C = 11π; A = 44π b. C = 22π; A = 44π c. C = 11π; A = 121π d. C = 22π; A = 121π

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

\[d=2r\]\[C=2\pi r\]\[A=\pi r^2\]

OpenStudy (kc_kennylau):

1. Find the radius 2. Use \(C=2\pi r\) to find the circumference 3. Use \(A=\pi r^2\) to find the area 4. Wait for me till I finish shower (other ppl will guide you)

OpenStudy (kc_kennylau):

cya :)

OpenStudy (anonymous):

what would the radius be

OpenStudy (haseeb96):

d=2r

OpenStudy (anonymous):

does anyone know what the radius would be

OpenStudy (haseeb96):

use it

OpenStudy (haseeb96):

r=11 inches

OpenStudy (anonymous):

oh thnx

OpenStudy (anonymous):

ok wait ill brb real quick

OpenStudy (anonymous):

ok im back

OpenStudy (haseeb96):

C=69.08 inches A=37.94inches ^2

OpenStudy (anonymous):

ok that's wat I got

OpenStudy (anonymous):

ok now what do we do

OpenStudy (haseeb96):

i means that r=11 inches since,C=2pir C=22 pi since, A=pi r^2 A=121 pi this is the correct answer i means , D is the correct answer

OpenStudy (haseeb96):

@prowrestler my method help u too much for understand...

OpenStudy (nirmalnema):

diameter of circle= 22inches therefore radius = 11 inches (d=2r) now circumference of circle= 2pi R =2*pi*11 that is 22pi and area = pi R^2 = pi *11* 11 = 121 pi

OpenStudy (anonymous):

So the answer is D?

OpenStudy (haseeb96):

yes

OpenStudy (anonymous):

thnx can you help on another please

OpenStudy (haseeb96):

yeah sure

OpenStudy (nirmalnema):

sure..